OpenStudy (anonymous):

write the partial fraction decomposition for the rational expression. (3x^2 - 7x - 2)/ (x^3 - x)

OpenStudy (anonymous):

\[\frac{a}{x}+\frac{b}{x+1}+\frac{c}{x-1}\] \[a(x-1)(x+1)+bx(x-1)+cx(x+1)=3x^2 - 7x - 2\] put \(x=0\) get \(a=-2\) right away

OpenStudy (anonymous):

oops i mean put \(x=0\) get \(a=2\) right away

OpenStudy (anonymous):

then replace \(x\) by 1 to get \(c\) and \(x\) by -1 to get \(b\)

OpenStudy (anonymous):

@satellite73 how do u know which numbers to pick?

OpenStudy (anonymous):

i pick the ones that make one of the factors zero, so it drops out

OpenStudy (anonymous):

let me write the steps

OpenStudy (anonymous):

before i begin, is this like ok \[a(x-1)(x+1)+bx(x-1)+cx(x+1)=3x^2 - 7x - 2\]?

OpenStudy (anonymous):

*line

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok now if you replace \(x\) by 0 then the \(b\) and \(c\) terms are gone you get \[a(0-1)(0+1)+b\times 0(0-1)+c\times 0(0+1)=3\times 0^2 - 7\times - 2\]

OpenStudy (anonymous):

and end up with \(-a=-2\) making \(a=2\) with some practice you can do it with your eyes (although i messed up the first time)

OpenStudy (anonymous):

the point is the \(b\) and \(c\) are gone if \(x=0\) so you can really just write \[a(0+1)(0-1)=-2\]\[-a=-2\]\[a=2\]

OpenStudy (anonymous):

then again start with \[a(x-1)(x+1)+bx(x-1)+cx(x+1)=3x^2 - 7x - 2\] this time make \(x=-1\) to get rid of the \(a\) and \(c\) terms

OpenStudy (anonymous):

you get \[b(-1)(-1-1)=3(-1)^2-7\times -1-2\] \[2b=8\] \[b=4\]

OpenStudy (anonymous):

and finally replace \(x\) by \(1\) to get rid of the \(a\) and \(b\) terms although once you know \(a=2\) and \(b=4\) you can find it easily, but this method is still snappy enough there are other methods as well, but this one is one of the easiest

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

yw